OpenStudy (anonymous):

Just a quick math question. Thanks! PS. im new. thanks again. What is the slope of the line 4x – 8y – 16 = 0 in simplest form? ____

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Samus_Aran

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Also, i will medal and fan.

OpenStudy (anonymous):

(Found the people i pinged through google (before i had a account here))

ganeshie8 (ganeshie8):

Heyy welcome to Openstudy :)

ganeshie8 (ganeshie8):

\[4x - 8y - 16 = 0 \]

OpenStudy (anonymous):

@Tres_Jolie

OpenStudy (anonymous):

Is it "Simple Math" or is there a Complex way to it?

OpenStudy (anonymous):

what grade are you in and is this algebra

ganeshie8 (ganeshie8):

It will be easy to eyeball the slope if you rearrange that equation to `slope-intercept form` : \[y = \color{red}{m}x+b\] \(\color{Red}{m}\) is the slope here

OpenStudy (anonymous):

Yes and grade 8

OpenStudy (anonymous):

part of a semester test i have to take/finish, And b = y-intercept. i know the y = mx+b part

OpenStudy (anonymous):

so all i need here is slope?

OpenStudy (anonymous):

its a fraction, by the way.....(Thanks for the help lol)

OpenStudy (anonymous):

ooohhhh ok

OpenStudy (anonymous):

let me see give me a min

ganeshie8 (ganeshie8):

Exactly! go ahead and write the equation in `y = mx + b` form

OpenStudy (anonymous):

@austin11111

OpenStudy (anonymous):

it always confuses me to write it though....will be a few minutes.

OpenStudy (anonymous):

y = 4x-16 ?

OpenStudy (anonymous):

did i do it correct? lol

OpenStudy (anonymous):

@ Master_Chief_117 write out the formula with the numbers you have like y=mx+b 8=m*4+b

OpenStudy (anonymous):

The way i got it, is y = 4x - 16, So, 4/1 at that point?

OpenStudy (anonymous):

so did you use the formula

OpenStudy (anonymous):

Quite confusing :/

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

how did you get y = 4x - 16 ?

OpenStudy (anonymous):

Its kind of hard to explain, Its just a way my father showed me how to do it......One moment

OpenStudy (anonymous):

Oh crap Moms calling...brb for a few

OpenStudy (anonymous):

I have to go :/ Can you get me a answer please? ill be back in about a hour hopefully @ganeshie8 @Tres_Jolie

ganeshie8 (ganeshie8):

take you time, tag me again when you're back il be around :)

OpenStudy (anonymous):

alright cool. Thanks! :)

ganeshie8 (ganeshie8):

4/1 is wrong btw

OpenStudy (anonymous):

@ganeshie8 Ugh....i had to go to the store and while i was gone my older brother, just for "kicks and giggles" puts random crap in....and Now i failed it....Ugh :(

OpenStudy (anonymous):

Next time ill remember to log out